OpenStudy (anonymous):

What are the steps for: lim x->∞ (√(x^2+4x+1)-x) Thanks

OpenStudy (maheshmeghwal9):

Is this the question? \[\color{red}{\lim_{x \rightarrow \infty}(\sqrt{(x^2+4x+1)}-x.)}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ive rationalized it but I am stuck on where to go from there

OpenStudy (anonymous):

\[4x+1/\sqrt{x^2+4x+1}+x\]

OpenStudy (maheshmeghwal9):

So carefully note that it is a type of question of limit At \[\pm \infty.\]with a radical in it. So If i do two examples for u . Then, would u be able to do ur question?

OpenStudy (maheshmeghwal9):

?

OpenStudy (anonymous):

solve away

OpenStudy (anonymous):

I was just confused about the plus/minus you had but I realized it was because of the radicand

OpenStudy (maheshmeghwal9):

so what do u want now examples or ur question's solution. former one is better.

OpenStudy (anonymous):

examples would be great. thanks

OpenStudy (maheshmeghwal9):

k! suppose if we take this as first example: - \[\color{red}{\lim_{x \rightarrow \infty}\frac{x+3}{\sqrt{x^2+4}}.}\]k:)

OpenStudy (anonymous):

alright

OpenStudy (maheshmeghwal9):

Now, If we notice the highest power of 'x' in the denominator; we see it is \[\sqrt{x^2}=x^1\] clear upto this?

OpenStudy (anonymous):

yea

OpenStudy (maheshmeghwal9):

now the next step would be as: - what to do of highest power of x is this. \[\lim_{x \rightarrow \infty}{\frac{\frac{1}{x}(x+3)}{\frac{1}{x}\sqrt{x^2+4}}}\] clear?

OpenStudy (anonymous):

yes

OpenStudy (maheshmeghwal9):

k! then u can write the above as \[\LARGE{\lim_{x \rightarrow \infty}{\frac{1+\frac{3}{x}}{\sqrt{\frac{1}{x^2}}\sqrt{x^2+4}}}}\]clear?

OpenStudy (anonymous):

yep

OpenStudy (maheshmeghwal9):

now it is nothing but this: - \[\lim_{x \rightarrow \infty}{\frac{1+\frac{3}{x}}{\sqrt{1+\frac{4}{x^2}}}}\]

OpenStudy (maheshmeghwal9):

clear?

OpenStudy (anonymous):

yes

OpenStudy (maheshmeghwal9):

then if u apply limit here i mean that If u put x->infinite here then u will see this\[\color{blue}{\frac{3}{x}=0}\] and\[\color{blue}{\frac{4}{x^2}=0.}\]

OpenStudy (maheshmeghwal9):

clear?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it becomes 1/1=1

OpenStudy (maheshmeghwal9):

then what u left with is this\[\large{\color{purple}{\frac{1}{\sqrt{1}}=1}}\]

OpenStudy (maheshmeghwal9):

ya u gt right.

OpenStudy (anonymous):

but what happens when you have no denominator like my example?

OpenStudy (maheshmeghwal9):

then u just make denominator by conjugate method. clear?

OpenStudy (maheshmeghwal9):

any doubt?

OpenStudy (anonymous):

yea Im afraid so. I understood the example you showed me. I had been taught that method already. But with the problem I posted earlier, I can't seem to apply it even after I used the conjugate and rationalize the function.

OpenStudy (maheshmeghwal9):

wait for upto 5 minutes. first I will try that. k!

OpenStudy (anonymous):

yes

OpenStudy (maheshmeghwal9):

is ur answer \[\color{green}{\infty?}\]

OpenStudy (anonymous):

no the correct answer is 2

OpenStudy (maheshmeghwal9):

what!

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=lim+x-%3E ∞+%28√%28x%5E2%2B4x%2B1%29-x%29

OpenStudy (anonymous):

i don't think that link works, but if you plug it into wolfram alpha, it will show you the answer and the steps, but I was trying to find a more concise way of doing the problem. Usually wolfram uses a cumbersome approach

OpenStudy (maheshmeghwal9):

ya:) If I do what I did in my copy would u be able to find my error?

OpenStudy (maheshmeghwal9):

only i say try!

OpenStudy (anonymous):

doubt it. I can't figure it out myself.

OpenStudy (maheshmeghwal9):

k! np:) u see just what i did?

OpenStudy (anonymous):

no i didn't

OpenStudy (maheshmeghwal9):

I gt after rationalising\[\lim_{x \rightarrow \infty}{\frac{\frac{1}{x}(4x+1)}{{\sqrt{\frac{1}{x^2}}(\sqrt{x^2+4x+1}+x).}}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ahhh I see it.

OpenStudy (anonymous):

I didn't see how you could cancel out 1/x and 1/x^2. I didn't take into account that the radicand reduced it to 1/x

OpenStudy (maheshmeghwal9):

& then just multiplied it whole \[\lim_{x \rightarrow \infty}{\frac{4+\frac{1}{x}}{\sqrt{1+\frac{4}{x}+\frac{1}{x^2}}+x}}\]& applied the limit & gt \[\frac{4+0}{\sqrt{1+0+0}+\infty}.\]\[\LARGE{\color{red}{\frac{4}{\infty}=\infty.}}\]

OpenStudy (maheshmeghwal9):

&